OpenStudy (anonymous):

A.) Find the slope of the graph of f(x) = 1 - e^x at the point where it crosses the x -axis B.) Find the Equation of the tangent line to the curve at this point

myininaya (myininaya):

So we need to find the x-intercepts if any then find f' and do f'(x-intercept)

OpenStudy (anonymous):

i think the answer to a would be 1. Am i correct with this? I set 1 - e^x = 0 and worked from there

myininaya (myininaya):

1=e^x => x=0 right? thats the x-intercept

myininaya (myininaya):

so now find f' f'(x)=-e^x and do f'(0)

OpenStudy (anonymous):

so that would be -1 then

myininaya (myininaya):

yep yep thats the slope :)

OpenStudy (anonymous):

then plug in original to find the y point?

myininaya (myininaya):

so y=mx+b y=-1x+b we know a point on the line(x=0,y=1-e^{0})=(0,0) the y-intercept is b=0 y=-x is the tangent line